# Why Larp?! A Synthesis Paper on Live Action Roleplay in Relation to HCI Research and Practice

### v1

@@ 4. Practitioner Overview of Larp, Technology, and Digital Aspects
This section focuses on information gathered on HCI relevant aspects of larp outside formal academic publications. The information comes from online information sources, drawing upon prior knowledge within the expert group. This overview does not strive to be all encompassing or unbiased, as it is based on the group’s knowledge. Still, within those limitations, the attempt is to represent larp as broadly as possible to give an understanding of at least some of what is out there. Larp in its essence can be seen as non technological, as it focuses on persons interacting in a physical space. But with the increased entanglement between the digital and physical, such aspects have become more and more integrated into the larp community, larp organizing processes and the larps themselves.
@@ 4.0.1. Technology for arranging larps
In the early days of larp, paper fanzines were often a way to spread information and exchange ideas. Since then the larp community has been quick to adapt to new technologies used to facilitate both co-creation and community building. For instance email-lists, online forums, homepages, blogs, Facebook groups and Instagram accounts. Using IT-competences within the community, several unique digital tools for organising larps have also been developed  [^16] Underworld, https://underworldlarp.com/managing-your-larp-technology/, for instance automatic scheduling systems for larp conventions  [^17] Prolog, https://prolog2022.spelkonvent.se/, booking sites to buy larp tickets, and automatic systems to write, sort and distribute player characters  [^18] Ensemble, https://ensemble.nu/ (see Figure LABEL:fig:ensemble). For Jubensha larps, special apps are often used for signups and also to find and pair participants (Xiong, 2022).
@@ 4.0.2. Technology in larps
Figure 2. Blodsband reloaded, https://www.bbreloaded.se/, was a humorous post-apocalyptic larp campaign that engaged with technology in multiple ways, for instance a digital system for communication, maps and hacking the in-game satellites, and the Radio Active recording studio, recording and broadcasting news and music to the whole larp. Here a player using the in-game hacking program at the larp. (Photo: Larpology)
There have also been a lot of different examples of technology used in larps themselves. These can be game mechanics such as steering and interacting with other spaceships  [^19] Voidship Concordia, https://spelkult.se/kalender/event/voidship-concordia/, chat functions to for instance chat with NPCs (non-player characters) who are off-site  [^20] Fight like a girl, https://flaglarp.wordpress.com/vision/, tech used for simulating (for instance magic), tracking player locations, aesthetic reasons and more  [^21] Mutanternas tidsålder, https://lajvverkstaden.se/. The tech can be diegetic or non-diegetic. Examples of technology and digital layers used are lights, soundscapes and smoke machines (Segura et al., 2017a). There are also examples of robotics being used, for instance an animatronic Velociraptor running around in the bunkers of the survival horror larp Insomnia: Imperium  [^22] Insomnia larp https://www.facebook.com/insomnialrp/. Jubensha larps utilise technology in many ways, for special effects and for instance to track player resources and stats (Xiong, 2022). Some examples of technology in larps are shown in Figures 2 and LABEL:fig:horizon.
There are organisations focusing specifically on developing technologies to be utilized at larps, such as Frivolous Engineering  [^23] https://www.facebook.com/FrivolousEngineering/, a non-profit organization based out of the Netherlands that designs technology for larps (Segura et al., 2017b). The technologies developed for larps can be used to create effects for immersion such as sound, lights and smoke, but they are also utilized to drive plots forward, contribute to character development, changing character agencies and steering the narrative.
A common solution at larps is that the technology is operated by Non-Player Characters (NPC), such as phones where you call a characters relatives, voiced by an NPC. Or, alternatively, larps use more advanced props for faking consciousness transfer (see Figure LABEL:fig:cmachine) from the larp Greylight  [^24] https://greylight.de/. But there are also technologies that are self-sustained and can be operated directly by the players, for example the fortune telling devices where characters received ominous messages about their future (see Figure LABEL:fig:vissers) at the larp Cirque Noir  [^25] https://nordiclarp.org/wiki/Cirque_Noir.
@@ 4.0.3. Fake technology in larps
Another way that larp incorporates technology is by creating and using fake technology. Larp’s immersive qualities and co-created storytelling make it possible to co-build a fiction that includes trying out technology that does not exist, and seeing how it affects persons, groups, lifestyles and society. For instance a common example of this is people playing androids (example in Figure LABEL:fig:mutant), or pretending to have been bio-hacked (example in Figure LABEL:fig:soldiersfightgirl) in different ways  [^26] Android Beginning, https://beginnings.atropos.se/. An example of this was the larp Last Will, a larp focusing on modern and future slavery. The larp was set in a dystopic future, where people in debt could become ”lifers” - that is slaves. Slaves got a necklace that the owners could remotely activate to deliver an electric shock, or just send out warnings of a shock. Shocks could be dealt individually, but also to the whole group of lifers, resulting in a collective punishment. This became a powerful game mechanic, despite it in reality just being simple led-light necklaces, and the ”remote controls” being ordinary whistles. It worked because everyone accepted and play-acted the fiction.
@@ 4.0.4. Larps about technology
There are many examples of larps exploring technology enhanced realities  [^27] https://alexandria.dk/en/data?scenarie=9900, such as dystopian, utopian, and what-if worlds, as in sci-fi, steampunk and post-apocalyptic genres.
Some larps have technological exploration as a main focus. For instance Omgiven av idioter (translated: Surrounded by idiots)282828‘Omgiven av idioter’, https://lajvkonvent.files.wordpress.com/2020/10/broschyr.pdf, in which participants lived in a near future where they lived all their lives via social media, and therefore needed special training in meeting people in real life (IRL). In this larp they were given devices that could, according to the fiction, show their emotions, color coded, as a wearable. Many larps engage with portraying and exploring different technological futures, and some larps dwell especially on philosophical and existential implications of technology. Examples of larps about technology are shown in Figures LABEL:fig:horizon, 4, 5, and 6.
Figure 4. Venue for the larp Zero - a sci-fi larp exploring future societies (Photo: Not Only Larp)
Figure 5. Turings Fråga was run in the Arch-Cathedral of Sweden, in collaboration with the Swedish Church. The larp was set in a non-defined future, and the participants were trying to pass a test to enter a prestigious academy. The test consisted of trying to identify the android in the group, and this was done using different criteria of what it means to be human (Photo: Johan Dahlberg)
@@ 4.0.5. Larps through technology
This type of experimentation with the format, and creating digitally hybrid larps, has been done using a wide range of different technological solutions. For instance there have been larps over discord (how, 2021), Zoom-larps (how, 2021), larps in VR  [^29] Ancient Hours, https://alexandria.dk/en/data?scenarie=14058 (VRl, 2023) and larps in Second Life  [^30] https://playabletheatre.org/project/slarp-fest-2021/. An online specific genre has even sprung up: online larp, or LAOG (Live Action Online Games) (Reininghaus, 2021), and there have also been academic courses on how to run online larps  [^31] https://lajvverkstaden.se/online-kurs-for-lajv-online-sodertorns-hogskola/. During Covid-19 Jubensha apps were launched and became very popular in China (Xiong, 2022).
While some of these concepts push the limits of what can be seen as larp, they clearly stem from the larp design tradition, and with roleplaying and social co-creative storytelling as a core. In some cases, for instance with larps over Zoom, the use of technology can be diegetic, meaning you dress up as and play a character that is sitting in front of a screen, so that this is a shared situation for both player and character (how, 2021). Examples of larps through technology are shown in Figures LABEL:fig:alchemist and LABEL:fig:ancient.
@@ 5. Literature Review: Larp in ACM
As mentioned, there are many larp-adjacent concepts in our field, but considering larp as a cultural phenomena and specific method, it is of interest to look at what has been done explicitly exploring larp in relation to HCI research. Thus, as part of the research process for this paper, a literature review was conducted in June 2023 using the ACM Digital Library. The inclusion criterion was that the term “larp” was used anywhere in the paper (body, title, keywords, etc.).”Larp” was used as the keyword in the search. All types of publications were included, and all years. For the purpose of the paper, it was important to narrow the scope of roleplaying to the cultural phenomenon of larp. Attempts were made to search also for related terms, such as roleplay, which have been used in many forms within the design and the games communities, but due to the wide use of the words roleplay, it was hard to untangle larp and larp-like experiences from non-larp, and therefore the inclusion criteria of explicitly using the word larp was settled upon. This first step resulted in 123 papers.
In a next step, all the articles were examined toward excluding irrelevant matches. The exclusion criterion (eliminating papers in which the abbreviation larp did not refer to live action roleplay), resulting in 71 papers remaining. After the selection process the abstracts and keywords for all the articles were read through, identifying HCI relevant concepts and terminologies mentioned. Themes were identified through a combination of keywords found in the articles, as well as based on the authors’ previous knowledge on topics related to larp.
The purpose of this literature review was to understand what topics and themes within HCI surrounding larp are frequently studied, in order to provide an overview of the field and identify future research directions and opportunities. In line with findings by Stefanindi et al. (Stefanidi et al., 2023), we state the purpose of the review and report our findings with regards to the purpose. Our review contains elements of both a systematic literature review (Moher et al., 2009) as well as a systematic mapping review (Epstein et al., 2020; Grant and Booth, 2009).
@@ 5.3. Common and Less Common Topics
Larp as a tool in design processes seems to be the most common type of paper about larp in our community. Interesting to note is that leisure larps are written about more often in the earlier years, but almost nothing is written about them in later years (Jonsson et al., 2006; Dagan et al., 2019a; Wohn et al., 2017; Vejdemo-Johansson et al., 2014; Segura et al., 2017a; Pichlmair et al., 2017; Dagan et al., 2019c; Oliveira et al., 2013; Márquez Segura et al., 2018). It is also worth noting that not a single paper, as found through this literature review, seems to focus on larps for formal education, despite this being a common type of edu-larp. There is a heavy slant towards focusing on adults and professionals. There is one article that focuses on children at a summer larp camp for girls (Fey et al., 2022).
Concepts such as pervasive games/pervasive play are more common in the earlier articles, while topics like wearables, design fiction and embodied interaction have been more common in later years. The research on larp seems to pick up on trends in the HCI community, utilising larps for different purposes and to explore current central topics. Most articles focus on designing larps, with the explicit aim of doing research. There seems to be a lack of articles looking at existing leisure larps, for instance doing interviews and observations of larpers and larp designers. One notable exception is the FDG’17 paper Design, Appropriation, and Use of Technology in Larps (Segura et al., 2017a), consisting of a survey of 39 larpers, investigating how they use digital technology in larps, finding around 300 examples of technology used, for different purposes and of different kinds.
@@ 5.4.2. Futuring
This common theme ties into the research traditions of design fiction (Dunne and Raby, 2013) and critical design (Bardzell et al., 2018). For instance, exploring the possibilities of wearables in relation to future tech and vulnerability. (Dagan et al., 2019d).
@@ 5.4.5. Playfulness
This is a recurring theme in many publications. Play has been a focus in HCI communities for a long time, even resulting in the creation of a specific venue for play-focused researchers, CHI PLAY  [^32] https://chiplay.acm.org/2023/. Regarding stances and theories within HCI, Ludic design (Frederique, 2021), as formulated by Gaver (Gaver et al., 2004), has been influential. There are elements of ludic design that seem to share characteristics with what has been done in relation to larp within the HCI community: playful ways of using design to explore concepts and sparking discussions. Drawing from, amongst other, examples of larp, the CHI’20 publication Technology for Situated and Emergent Play: A Bridging Concept and Design Agenda (Altarriba Bertran et al., 2020) specifically proposes a design agenda for the types of situated and emergent play that occur in larp. Exploring play and playfulness ties into research traditions much older than HCI, tapping into fields such as sociology and education, with for instance Huizinga (Huizinga, 1955), Piaget (Piaget, 2013), Vygotsky (Vygotsky, 1967) and Caillois (Caillois, 2001) as key references. Earlier work is often only very briefly touched upon in the publications, if at all. The Future Play’07 paper Pervasive Games in Ludic Society focuses especially on concepts of play and ludic, and looks at long-standing cultural trends influencing larp among other types of play, even though the paper focuses less on foundations in research traditions and looks more outside academia.
@@ 5.4.6. Appropriation, co-creation and Participatory Design
This is another set of theories that has been identified in the publications surveyed. This ties into for instance Harrison’s and Dourish’s theories on space and place (Harrison and Dourish, 1996), and also for instance Designing for Appropriation (Dix, 2007). A publication taking this as a focus is Design, appropriation and use of technology in larps (Segura et al., 2017a), presented at FDG ’17. The TEI’21 publication Designing Gaming Wearables: From Participatory Design to Concept Creation (Jung et al., 2021) is an example of a paper that explicitly focuses on larp and participatory design.
@@ 5.5. Limitations
As was mentioned, this overview is limited to only those articles in the ACM database explicitly mentioning larp. There are probably also more articles that uses larp-like methods without using the term. Further, there are many publications outside ACM, especially within the field of game studies  [^33] The Transformative Play Initiative, https://www.speldesign.uu.se/research/games-and-society-lab/transformative-play/Transformative-Play-Initiative/and education (Geneuss, 2019), and there is even peer-reviewed journals; the International Journal of Role-playing  [^34] https://journals.uu.se/IJRP/index, and Japanese Journal of Analog Role-Playing Game Studies  [^35] https://jarps.net/journal,focusing on roleplays such as larp. This paper focuses specifically on larp in an HCI context, not on exploring larp or roleplay as a broader phenomenon. The literature review was conducted focusing on titles, abstracts and keywords. Statistics on affiliations and authors are based on ACMs meta-data. Overlapping authors who have changed affiliations etc. have not been tracked.

### v2

@@ 4. Practitioner Overview of Larp, Technology, and Digital Aspects
This section focuses on information gathered on HCI relevant aspects of larp outside formal academic publications. The information comes from online information sources, drawing upon prior knowledge within the expert group. This overview does not strive to be all encompassing or unbiased, as it is based on the group’s knowledge. Still, within those limitations, the attempt is to represent larp as broadly as possible to give an understanding of at least some of what is out there. Larp in its essence can be seen as non technological, as it focuses on persons interacting in a physical space. But with the increased entanglement between the digital and physical, digital aspects have become more and more integrated into the larp community, larp organizing processes and the larps themselves.
@@ 4.0.1. Technology for arranging larps
In the early days of larp, paper fanzines were often a way to spread information and exchange ideas. Since then, the larp community has been quick to adapt to new technologies used to facilitate both co-creation and community building, using email-lists, online forums, homepages, blogs, Facebook groups and Instagram accounts. Using IT-competences within the community, several unique digital tools for organising larps have also been developed  [^10] Underworld, https://underworldlarp.com/managing-your-larp-technology/, for instance automatic scheduling systems for larp conventions  [^11] Prolog, a swedish larp convention https://prolog2022.spelkonvent.se/, booking sites to buy larp tickets, and automatic systems to write, sort and distribute player characters  [^12] Ensemble, https://ensemble.nu/ (see Figure LABEL:fig:ensemble). For Jubensha larps, special apps are often used for signups and also to find and pair participants (Xiong, 2022).
@@ 4.0.2. Technology in larps
Figure 2. Blodsband reloaded larp. Here is a player using the in-game hacking program at the larp (see 7 in Appendix A). (Photo: Larpology)
There have also been a lot of different examples of technology used in larps themselves. These can be game mechanics such as steering and interacting with other spaceships (see 22 in Appendix A), chat functions to for instance chat with NPCs (non-player characters) who are off-site (see 8 in Appendix A), tech used for simulating (for instance magic), tracking player locations, aesthetic reasons and more (see 14 in Appendix A). The tech can be diegetic (embedded into the larp narrative) or non-diegetic (in the background). Examples of technology and digital layers used are lights, soundscapes and smoke machines (Segura et al., 2017). There are also examples of robotics being used, for instance an animatronic Velociraptor running around in the bunkers of the survival horror larp Insomnia: Imperium (see 12 in Appendix A. Jubensha larps utilise technology in many ways, for special effects and for instance to track player resources and stats (Xiong, 2022). Some examples of technology in larps are shown in Figures 2 and LABEL:fig:horizon. Utilizing technologies at larps turns them into a type of hybrid experience, that changes the way we should think about designing the experience. Hybrid experiences, which mix physical, digital and technological components requires special design considerations, as has been seen in for instance research on hybrid games (Sparrow and Rogerson, 2023).
There are organisations focusing specifically on developing technologies to be utilized at larps, such as Frivolous Engineering  [^13] Frivolous Engineering, https://www.facebook.com/FrivolousEngineering/, a non-profit organization based out of the Netherlands that designs technology for larps (Segura et al., 2017). The technologies developed for larps can be used to create effects for immersion such as sound, lights and smoke, but they are also utilized to drive plots forward, contribute to character development, changing character agencies and steering the narrative.
A common solution at larps is that the technology is operated by Non-Player Characters (NPC), such as phones where you call a characters relatives, voiced by an NPC. Or, alternatively, larps use more advanced props such as faked medical facilities (see Figure LABEL:fig:cmachine from the larp Mission Together), a machine for faking consciousness transfer at the larp Greylight 2142 (see 10 in Appendix A), and the control stations at the larp Odysseus (see Figure LABEL:fig:vissers and further detailed in entry 26 of Appendix A). But there are also technologies that are self-sustained and can be operated directly by the players, for example the fortune telling device made by the larp prop company Frivolous Engineering, where characters received ominous messages about their future, used for instance at the larp Cirque Noir  [^14] Cirque Noir https://nordiclarp.org/wiki/Cirque_Noir.
@@ 4.0.3. Fake technology in larps
Figure 4. Shows Mutanternas tidsålder larp, (Photo: LajvVerkstaden), see 14 in Appendix A
Another way that larp incorporates technology is by creating and using fake technology. Larp’s immersive qualities and co-created storytelling make it possible to co-build a fiction that includes trying out technology that does not exist, and seeing how it affects persons, groups, lifestyles and society. This kind of fake technologies and simulations bear resemblance to design methods such as Wizard of Oz (Kelley, 2018) and rapid prototyping. A common example of fake technology at larps is people playing androids (example in Figure 4), or pretending to have been bio-hacked (see 3 in Appendix A). An example of this was the larp Last Will (see 24 in Appendix A), a larp focusing on modern and future slavery. The larp was set in a dystopic future, where people in debt could become “lifers” - that is slaves. Slaves got a necklace that the owners could remotely activate to deliver an electric shock, or just send out warnings of a shock. Shocks could be dealt individually, but also to the whole group of lifers, resulting in a collective punishment. This became a powerful game mechanic, despite it in reality just being simple led-light necklaces, and the “remote controls” being ordinary whistles. It worked because everyone accepted and play-acted the fiction.
@@ 4.0.4. Larps about technology
There are many examples of larps exploring technology enhanced realities (see 6 in Appendix A), such as dystopian, utopian, and what-if worlds, as in sci-fi, steampunk and post-apocalyptic genres.
Some larps have technological exploration as a main focus. For instance Omgiven av idioter (see 17 in Appendix A, translated to ‘Surrounded by idiots’, in which participants lived in a near future where they lived all their lives via social media, and therefore needed special training in meeting people in real life (IRL). In this larp they were given devices that could, according to the fiction, show their emotions, color coded, as a wearable. Many larps engage with portraying and exploring different technological futures, and some larps dwell especially on philosophical and existential implications of technology. Examples of larps about technology are shown in Figures LABEL:fig:horizon, 6, 7, and 8.
Figure 6. Venue for the larp Mission Together - a sci-fi larp exploring future societies, arranged by Not Only Larp (Photo: Kai Simon Fredriksen). See 26 in Appendix A for more information about the larp.
Figure 7. Turings Fråga (see 21 in Appendix A) was run in the Arch-Cathedral of Sweden, in collaboration with the Swedish Church, organized by Dahlberg, Engström and Gamero. The larp was set in a non-defined future, and the participants were trying to pass a test to enter a prestigious academy. The test consisted of trying to identify the android in the group, and this was done using different criteria of what it means to be human (Photo: Johan Dahlberg)
@@ 4.0.5. Larps through technology
This type of experimentation with the format, and creating digitally hybrid larps, has been done using a wide range of different technological solutions. For instance there have been larps over discord (how, 2021), Zoom-larps (how, 2021), larps in VR (see 2 in Appendix A) and larps in Second Life  [^15] Playable Theatre, https://www.indiecade.com/slarp-fest-2021/. An online specific genre has even sprung up: online larp, or LAOG (Live Action Online Games) (Reininghaus, 2021), and there have also been academic courses on how to run online larps 161616 LajvVerkstaden, https://lajvverkstaden.se/online-kurs-for-lajv-online-sodertorns-hogskola/. During Covid-19 Jubensha apps were launched and became very popular in China (Xiong, 2022).
While some of these concepts push the limits of what can be seen as larp, they clearly stem from the larp design tradition, and with roleplaying and social co-creative storytelling as a core. In some cases, for instance with larps over Zoom, the use of technology can be diegetic, meaning you dress up as and play a character that is sitting in front of a screen, so that this is a shared situation for both player and character (how, 2021). An example of larps through technology is shown in Figure 9, as well as in the larp Ancient Hours, detailed further in entry 2 of Appendix A.
Figure 9. The alchemists classroom (note the microphone in the front). (Photo by LajvVerkstaden).
@@ 5. Literature Review: Larp in ACM
As mentioned, there are many larp-adjacent concepts in our field, but considering larp as a cultural phenomena and specific method, it is of interest to look at what has been done explicitly exploring larp in relation to HCI research. Thus, as part of the research process for this paper, a literature review was conducted in June 2023 using the ACM Digital Library. The inclusion criterion was that the term “larp” was used anywhere in the paper (body, title, keywords, etc.). “Larp” was used as the keyword in the search. All types of publications were included, and all years. For the purpose of the paper, it was important to narrow the scope of roleplaying to the cultural phenomenon of larp. Attempts were made to search also for related terms, such as roleplay, which have been used in many forms within the design and the games communities, but due to the wide use of the words roleplay, it was hard to untangle larp and larp-like experiences from non-larp, and therefore the inclusion criteria of explicitly using the word larp was settled upon. This first step resulted in 123 papers.
In a next step, all the articles were examined toward excluding irrelevant matches. The exclusion criterion (eliminating papers in which the abbreviation larp did not refer to live action roleplay), resulting in 71 papers remaining. After the selection process the abstracts and keywords for all the articles were read through, identifying HCI relevant concepts and terminologies mentioned. Themes were identified through a combination of keywords found in the articles, as well as based on the authors’ previous knowledge on topics related to larp, thus it is a combination of bottom-up and top-down thematic analysis.
The purpose of this literature review was to understand what topics and themes within HCI surrounding larp are frequently studied, in order to provide an overview of the field and identify future research directions and opportunities. Our review contains elements of both a systematic literature review (Moher et al., 2009) as well as a systematic mapping review (Epstein et al., 2020; Grant and Booth, 2009).
@@ 5.3. Common and Less Common Topics
Larp as a tool in design processes seems to be the most common type of paper about larp in our community. Interesting to note is that leisure larps are written about more often in the earlier years, but almost nothing is written about them in later years (Jonsson et al., 2006; Dagan et al., 2019a; Wohn et al., 2017; Vejdemo-Johansson et al., 2014; Segura et al., 2017; Pichlmair et al., 2017; Dagan et al., 2019c; Oliveira et al., 2013; Márquez Segura et al., 2018). It is also worth noting that not a single paper, as found through this literature review, seems to focus on larps for formal education, despite this being a common type of edu-larp. There is a heavy slant towards focusing on adults and professionals. There is one article that focuses on children at a summer larp camp for girls (Fey et al., 2022).
Concepts such as pervasive games/pervasive play are more common in the earlier articles, while topics like wearables, design fiction and embodied interaction have been more common in later years. The research on larp seems to pick up on trends in the HCI community, utilising larps for different purposes and to explore current central topics. Most articles focus on designing larps, with the explicit aim of doing research. There seems to be a lack of articles looking at existing leisure larps, for instance doing interviews and observations of larpers and larp designers. One notable exception is the FDG’17 paper Design, Appropriation, and Use of Technology in Larps (Segura et al., 2017), consisting of a survey of 39 larpers, investigating how they use digital technology in larps, finding around 300 examples of technology used, for different purposes and of different kinds.
@@ 5.4.2. Futuring
This common theme ties into the research traditions of design fiction (Dunne and Raby, 2013) and critical design (Bardzell et al., 2018). For instance, exploring the possibilities of wearables in relation to future tech and vulnerability (Dagan et al., 2019d).
@@ 5.4.5. Playfulness
This is a recurring theme in many publications. Play has been a focus in HCI communities for a long time, even resulting in the creation of a specific venue for play-focused researchers, CHI PLAY  [^17] CHI PLAY 2023 website, https://chiplay.acm.org/2023/. Regarding stances and theories within HCI, Ludic design (Frederique, 2021), as formulated by Gaver (Gaver et al., 2004), has been influential. There are elements of ludic design that seem to share characteristics with what has been done in relation to larp within the HCI community: playful ways of using design to explore concepts and sparking discussions. Drawing from, amongst other, examples of larp, the CHI’20 publication Technology for Situated and Emergent Play: A Bridging Concept and Design Agenda (Altarriba Bertran et al., 2020) specifically proposes a design agenda for the types of situated and emergent play that occur in larp. Exploring play and playfulness ties into research traditions much older than HCI, tapping into fields such as sociology and education, with for instance Huizinga (Huizinga, 1955), Piaget (Piaget, 2013), Vygotsky (Vygotsky, 1967) and Caillois (Caillois, 2001) as key references. Earlier work is often only very briefly touched upon in the publications, if at all. The Future Play’07 paper Pervasive Games in Ludic Society (Stenros et al., 2007) focuses especially on concepts of play and ludic, and looks at long-standing cultural trends influencing larp among other types of play, even though the paper focuses less on foundations in research traditions and looks more outside academia.
@@ 5.4.6. Appropriation, co-creation and Participatory Design
This is another set of theories that has been identified in the publications surveyed. This ties into for instance Harrison and Dourish’s theories on space and place (Harrison and Dourish, 1996), and also for instance Designing for Appropriation (Dix, 2007). A publication taking this as a focus is Design, appropriation and use of technology in larps (Segura et al., 2017), presented at FDG ’17. The TEI’21 publication Designing Gaming Wearables: From Participatory Design to Concept Creation (Jung et al., 2021) is an example of a paper that explicitly focuses on larp and participatory design.
@@ 5.5. Limitations
As was mentioned, this overview is limited to only those articles in the ACM database explicitly mentioning larp. There are probably also more articles that uses larp-like methods without using the term. Further, there are many publications outside ACM, especially within the field of game studies  [^18] The Transformative Play Initiative, https://www.speldesign.uu.se/research/games-and-society-lab/transformative-play/Transformative-Play-Initiative/ and education (Geneuss, 2019), and there is even peer-reviewed journals; the International Journal of Role-playing  [^19] International Journal of Role-playing, https://journals.uu.se/IJRP/index, and Japanese Journal of Analog Role-Playing Game Studies  [^20] Japanese Journal of Analog Role-playing Game Studies, https://jarps.net/journal,focusing on roleplays such as larp. This paper focuses specifically on larp in an HCI context, not on exploring larp or roleplay as a broader phenomenon. The literature review was conducted focusing on titles, abstracts and keywords. Statistics on affiliations and authors are based on ACMs meta-data. Overlapping authors who have changed affiliations etc. have not been tracked.